# Joint Liability Model with Adaptation to Climate Change

### v1

@@ Abstract
This paper extends the application of ESG score assessment methodologies from large corporations to individual farmers’ production, within the context of climate change. Our proposal involves the integration of crucial agricultural sustainability variables into conventional personal credit evaluation frameworks, culminating in the formulation of a holistic sustainable credit rating referred to as the Environmental, Social, Economics (ESE) score. This ESE score is integrated into theoretical joint liability models, to gain valuable insights into optimal group sizes and individual-ESE score relationships. Additionally, we adopt a mean-variance utility function for farmers to effectively capture the risk associated with anticipated profits. Through a set of simulation exercises, the paper investigates the implications of incorporating ESE scores into credit evaluation systems, offering a nuanced comprehension of the repercussions under various climatic conditions.
@@ 1 Introduction
1.1 Credit Scoring System
Recently, there has been a growing interest worldwide in improving agricultural productivity, driven by concerns mainly over increasing food insecurity and population growth, in the face of climate change. One potential solution to these challenges is to provide better access to institutional credit, which includes microcredit, commercial banks, and agricultural banks, as pointed out by Römer and Mußhoff (2018). This improved access can help rural households to mitigate risks and obtain essential resources and technologies to modernize agricultural practices, as well as facilitate the development of farm/nonfarm linkages. Khandker and Koolwal (2016), for instance, argue that institutional credit enables rural households to achieve better access to technology and inputs, which can result in a modernized agricultural system. By reducing risk, institutional credit can improve the overall resilience of rural households, while simultaneously bolstering the rural economy. Through this approach, rural households can leverage institutional credit to access new markets and improve their overall productivity, creating thereby a more sustainable agricultural system that is better equipped to meet the demands of a fastly changing world in the face of global climate change.
Microfinance credit scoring has traditionally been found to be an effective tool for improving risk management in the microfinance sector. The use of credit scoring models has been shown to increase the efficiency of loan assessment in Bumacov et al. (2014), leading to increased productivity of loan officers. Additionally, Schreiner (2009) highlighted that credit scoring models are objectively able to reflect complex causal relationships, which improves the accuracy of credit risk assessments. The automated nature of credit scoring further assists loan officers in making lending decisions, leading to a reduction in credit risk and loan default. The use of credit scoring also leads to cost reductions in the microfinance sector. Van Gool et al. (2012) recommended incorporating credit scoring into the lending process as a means to integrate statistical and human best practices, thereby enhancing the precision of lending decisions. This claim is supported by evidence assembled in the credit-scoring literature that exists for developing countries in various regions of the world including Africa, Asia, Eastern Europe, and Latin America, as mentioned in Römer and Mußhoff (2018). In most of these regions, credit scoring models have been found to be suitable for determining credit risk or supporting loan officers in their decision-making processes.
1.2 Joint Liability Contracts
Joint liability is a common feature of many lending schemes, especially in the context of microfinance. In a joint liability scheme, multiple borrowers are held collectively responsible for repaying a loan. This can have important implications for borrowers, lenders, and even the broader communities. Joint liability has been shown to be effective in promoting repayment rates and in reducing default risk. For example, a study by Ghatak and Guinnane (1999) find that joint liability groups had significantly lower default rates than individual borrowers. Similarly, Khandker et al. (1998) find that group-based lending in Bangladesh increased the number of loans made to poor borrowers and reduced the cost of delivering credit. Microfinance has gained widespread recognition as a powerful instrument for poverty reduction and technology adoption by offering credit access to underprivileged households. Group lending is among the factors that have contributed to microfinance’s success. As mentioned earlier, joint liability group lending has been promoted as a means of reducing default rates and enhancing community welfare by addressing four important drawbacks of conventional individual lending: adverse selection, moral hazard, auditing expenses, and enforcement of Loan Terms. Nevertheless, recent studies have raised doubts about the superiority of joint liability group lending over individual lending in terms of its impacts on borrowers and lenders. Madajewicz (2011), for example, compares joint liability group lending to individual lending in the presence of heterogeneity among borrowers. The authors show that joint liability group lending can be beneficial in terms of reducing adverse selection and moral hazard, but it may not be superior to individual lending in terms of efficiency and borrower welfare. Group loans may allow for larger loans for the poorest borrowers, but individual loans may be more beneficial for wealthier, credit-constrained borrowers. For these borrowers, individual loans may lead to larger businesses than group loans. In addition, the findings in Rezaei et al. (2017) show that group size can affect lending efficiency. Larger groups are more reliable in repaying loans for riskier projects, but the group size cannot become too large. Joint liability has a positive effect on borrowers’ repayment amount and welfare compared to individual liability, provided that group members impose strong punishment strategies against any strategically defaulting member. Rezaei et al. (2017) also examined the possibility of relaxing some of the simplifying assumptions made in the basic model of joint liability. Under the climate change context, Castaing (2021) investigates the impact of risk-sharing networks on the adoption of weather shock management strategies in Burkina Faso; particularly it studies how the principle of forced solidarity may reduce efforts to adopt techniques that mitigate exposure to climate change. The study finds that a system based on mutual assistance between farmers may reduce their incentives to adopt risk-mitigating strategies, both incrementally and transformationally. The study highlights the need for complementary mechanisms, such as insurance, to help farmers reduce their vulnerability to climate change.
As the world becomes increasingly aware of the impact of climate change on the agricultural sector, there is a growing emphasis on the importance of sustainable agricultural practices. In particular, the concept of green agriculture and sustainable development has gained considerable attention in recent times with the development of agricultural technology. This emphasis on sustainability is particularly relevant in the context of agricultural projects, where the potential for sustainable practices can have a major impact on the future of these projects.
1.3 Sustainable Scoring System
Cumming et al. (2017) emphasized that sustainable agriculture practices contribute to the achievement of the Sustainable Development Goals (SDGs) set by the United Nations, which aim at addressing global challenges such as poverty, hunger, climate change, and environmental degradation. For example, sustainable agriculture practices help to conserve natural resources such as water, soil, and biodiversity, which are critical for human survival and well-being. Sustainable farming also helps to reduce greenhouse gas emissions and mitigate climate change by sequestering carbon in the soil and reducing the use of synthetic fertilizers and pesticides that contribute to emissions. Moreover, sustainable agriculture practices can help farmers increase their productivity, reduce their costs, and enhance their resilience to climate change, market volatility, and other risks. Sustainable farming practices can also improve the quality and safety of the food produced, reducing food insecurity and malnutrition in local and global populations, as argued by Calicioglu et al. (2019).
There is a growing interest among academics and practitioners to integrate the ESG score in the financial decision-making process to evaluate the risk profile of large companies. This trend has inevitably also been extended to the agricultural sector, which calls for sustainable agricultural practices and the use of sustainable metrics to evaluate agricultural activities. In their report, Rivera-Ferre (2008) discussed the future of the agriculture sector in the economy and provided a definition for sustainable agricultural practices. According to this report, sustainable agricultural practices are those that are environmentally, socially, and economically sustainable. Firstly, from an environmental standpoint, it avoids inefficient use of water resources, loss of biodiversity due to converting natural habitat into agricultural land, imprudent use of chemical compounds that negatively affect local and regional water and air quality, emissions of greenhouse gases that disrupt the global climate, and losses in soil health and fertility. Secondly, from an economic perspective, sustainable agriculture enhances the economic viability of the agricultural sector by improving agricultural productivity and profitability, advancing agricultural innovation, facilitating farmers’ access to markets and credit, improving farmers’ ability to manage risk, and reducing food losses throughout the supply chain. Lastly, from a social perspective, sustainable agriculture enhances farmers’ well-being, respects farmers’ rights, promotes equitable opportunities in rural communities, and benefits all of society by improving food supply system resilience and enhancing nutrition and health. Zhang et al. (2021) discussed the challenges of defining and measuring sustainable agriculture and the need for independent and transparent measurements of national sustainability in this field. They developed a Sustainable Agriculture Matrix (SAM) to quantify national performance indicators in agriculture and investigated the trade-offs and synergies based on historical data for most countries of the world. The SAM provides a useful starting point for evaluating progress, identifying priorities for improvement, and informing national policies and actions toward sustainable agriculture.
@@ 1.3 Sustainable Scoring System
In recent years, the importance of incorporating sustainability factors into financial decision-making has gained traction. While credit metrics for corporations have been extensively studied, there remains a noticeable gap in the assessment of individual borrowers, particularly farmers seeking credit to finance their agricultural operations. To address this gap, this paper introduces a comprehensive scoring system known as an Environmental, Sustainability, and Economics (ESE), which considers not only traditional credit metrics, including economic factors, but importantly also environmental and social aspects that contribute to sustainable agricultural practices.
The integration of sustainability factors into credit evaluation processes is vital for both lenders and borrowers in the agricultural sector. For lenders, the ESE framework enables the assessment of a borrower’s long-term financial viability, identification of risks and opportunities associated with sustainability practices, and ensures alignment with ESE principles. From the borrower’s perspective, it demonstrates a commitment to sustainability, improves access to capital, and provides opportunities for cost savings through energy and resource efficiency. The ESE framework is particularly relevant for individual farmers, as it acknowledges the importance of balancing environmental and sustainability goals with economic viability. Farmers are more likely to adopt sustainable practices if they can see clear economic benefits from doing so. In the agricultural sector, adopting sustainable farming practices often requires initial investments, such as new equipment, technology, or training. However, if these practices lead to cost savings, increased productivity, or other economic advantages in the long run, farmers will be more inclined to adopt them.
Furthermore, the ESE framework helps lenders better assess the long-term financial viability of farmers who adopt sustainable practices. By considering the potential economic benefits of environmentally friendly and socially responsible farming practices, lenders can make more informed decisions about extending credit to borrowers, ultimately encouraging the adoption of sustainable practices across the agricultural sector. In summary, this paper stresses the critical link between environmental sustainability and economic viability for individual farmers through the ESE framework. This comprehensive approach ensures that sustainable practices are not only environmentally and socially beneficial, but also economically viable, which is essential to promote widespread adaptation measures and to sustain a long-term success.
Monitoring, reporting, and verification (MRV) is a system used to track and assess the effectiveness of actions taken to address environmental or social issues. It involves the collection and analysis of data to monitor progress toward specific goals or targets, reporting of that information to stakeholders, and verification of the results by an independent third party to ensure accuracy and credibility. The purpose of an MRV system is to provide transparent and reliable information that can be used to guide decision-making, assess the effectiveness of interventions, and demonstrate accountability to stakeholders. Examples of areas where MRV systems are commonly used include climate change mitigation, biodiversity conservation, and social impact assessment. MRV systems can be applied at different scales, from individual projects to national or international initiatives. While the MRV system and the Individual-ESE Score Mechanism appear to be two separately distinct tools, they are actually interconnected with each other. Specifically, the data collected through the MRV system can be used to update and refine the ESG scoring system, ensuring that it remains relevant and accurate over time. For example, if the MRV system shows that a particular Individual-ESE factor has a more significant impact than previously thought, the Individual-ESE Score Mechanism can be adjusted to reflect this new piece of information. By using data from the MRV system to update the Individual-ESE scoring system, lenders and borrowers can therefore make better-informed decisions and ensure that their joint liability contract is sustainable over a long-term period. Therefore, the MRV system and the Individual-ESE Score Mechanism can complement each other well and together they can enhance the effectiveness of joint liability contracts. Tang et al. (2018) listed limitations of the MRV system, including ambiguity in the legal status of relevant policies and regulations, unclear requirements for monitoring plans, the lack of consistency and harmonization in accounting and reporting guidelines, and the lack of information technology (IT); This calls for further research in this area.
1.4 Motivation and Outline
@@
The motivation for this paper is to holistically evaluate a more comprehensive profile of individual farmer borrowers and provide a more accurate assessment of their creditworthiness. Creating an individual-ESE score for individual borrowers in a joint liability problem can help to create more sustainable joint liability contracts. The resulting individual-ESE score would be integrated into a joint liability model as an additional criterion for assessing creditworthiness. In this paper, we take a first but critical step in performing a theoretical exercise by linking the Individual-ESE score to our theoretical models and providing lenders and borrowers with optimal scores and corresponding optimal group sizes by maximizing the total utility of a group of borrowers. We start our analysis with the simplest case, which only contains two borrowers in the group. Later we generalize it to a group of $n$ individual borrowers. It is important to incorporate the Individual-ESE Score Mechanism into our joint liability model, which can also benefit both lenders and borrowers. For lenders, this mechanism provides a better understanding of the borrower’s risk profile, enabling them to make informed lending decisions and, in the process, ensure more sustainable joint liability contracts. It also provides information about the optimal group size for lending. For borrowers, the mechanism provides a clearer picture of their ranking relative to their peers, allowing them to identify strengths and weaknesses and work on improving ESE scores. It also provides information about the number of peers needed for a group borrowing purpose, helping borrowers form groups that are more likely to succeed and have improved access to credit. Incorporating the Individual-ESE Score Mechanism into credit evaluation processes can thus lead to better-informed lending decisions, improved access to credit, and sustainable development outcomes.
@@
In the remaining parts of this paper, we list the steps and potential metrics for building an individual-ESE score mechanism in Section 2. Next, we introduce the model setup and the assumptions in Section 3, and derive a set of results from proposed theoretical models of a group size of two and then a group size of $n$ in Section 4 and Section 5 respectively. In Section 6, we refine the utility function to a mean-variance form, which incorporates both the expected value of the return and risk simultaneously in the face of climate change. The resulting model does not have closed-form analytical expressions and, as a result, a set of simulation exercises is provided through a set of climate scenarios.
2 The Mechanism of Individual-ESE Scores
In order to uphold accountability for sustainable agriculture commitments and provide valuable insights for policymaking, conducting regular and transparent assessments is of paramount importance. However, there are a number of variations in the definitions of sustainable agriculture, and also a limited number of quantitative assessments of agricultural sustainability that currently are available, according to Agovino et al. (2018). While some experts view sustainable agriculture as a collection of management strategies, others see it as an ideology or a specific set of goals. Despite this, there is an increasing consensus that sustainable agriculture should be evaluated based on its environmental, economic, and social impacts, as emphasized by Zhang et al. (2021). Several frameworks and indicators have been developed. Existing sustainable agriculture indicators mainly assess the environmental impacts of agriculture production, and many have limited data availability. Although various organizations have called for monitoring agriculture globally, actual datasets to enable trend assessments are still currently lacking. The absence of a consistent quantification of agricultural sustainability across multiple dimensions impedes the successful identification of undesirable trade-offs and the development of win-win solutions. Given these challenges, relying on literature-based metrics is a reasonable starting point in developing a sustainable agricultural credit scoring system. Literature-based metrics are metrics that have been previously used and validated in other studies or publications and are therefore based on a larger sample size and a more diverse set of variables than any single data collection effort. Having literature-based metrics about agricultural credit scoring systems allows us to apply them when we obtain real data or use them to create a scoring system by using synthetic data. By compiling these synthetic metrics and adjusting them to fit the specific needs of the agricultural sector, we can create a credit scoring system that is more comprehensive and accurate than any single data collection effort.
The ultimate objective of this research is to eventually establish a comprehensive and sustainable agricultural credit scoring system at the individual level, which encompasses three critical aspects: Economic, Social, and Environmental. This system will be called an individual-ESE score. While traditional credit scoring systems focus mainly on borrowers’ economic and financial status, our system will take into account a broader range of factors that contribute to sustainability. We recognize the importance of economic metrics and acknowledge that they are crucial to the classical credit scoring system. Therefore, we will use some of the commonly used economic metrics in the related literature as a foundation for our economic metrics. Presented below are synopses of selected academic papers that have conducted research on the agricultural credit scoring system. (1) Römer and Mußhoff (2018) investigated the use of weather data in agricultural credit scoring for microfinance institutions, finding that simple correlation analysis between weather events and loan repayment is insufficient for forecasting future repayment behavior. (2) Turvey (1991) reviewed the use of credit scoring for agricultural loans and its application to agricultural finance, identifying the importance of accurate data and the need for specialized credit scoring models for agricultural lending. (3) Lufburrow et al. (1984) developed a credit scoring model for farm loan pricing, utilizing statistical methods to determine the significant factors affecting loan repayment and default rates. (4) Limsombunchai et al. (2005) analyzed credit scoring for agricultural loans in Thailand, identifying the factors that affect credit scores and their impact on loan approval and interest rates. (5) Bumacov et al. (2014) examined the use of credit scoring in microfinance institutions, reporting that it can improve the outreach and efficiency of microfinance operations, but also highlighting the need for effective risk management strategies. (6) Mutamuliza et al. (2021) investigated the determinants of smallholder farmers’ participation in microfinance markets in Rwanda, identifying factors such as gender, education, and access to information that can affect farmers’ ability to access credit. We provide a summary of the key metrics employed in their credit scoring models as the foundations of economic metrics in Table 1.
Zhang et al. (2021) presents a variety of useful metrics for the environmental and social pillars. With additional information provided in Fuglie et al. (2016) and Movilla-Pateiro et al. (2021), we have compiled a comprehensive list of potential metrics for these pillars, along with detailed descriptions and methods of measurement. The metrics are shown in Table 2.
Metric Description Used as a Metric in the Scoring System
Age Age of the applicant in years (1), (4)
Gender Male, Female, Others (1), (6)
Education Education level of the applicant (6)
Collateral Collateral in thousands (1), (3), (4)
Assets Assets in thousands (1), (2), (3), (4), (5)
Debt Debt to other banks in thousands (1), (2), (3), (4), (5)
Deposit Deposit in the bank account in thousands (1), (2), (5)
Income Monthly business and household income in thousands (1), (2), (3), (4), (6)
Marital status Single, Married, Divorced, Others (1)
No. family members Number of family members (1)
Loan Size Loan amount in thousands (1), (4), (6)
Loan duration Loan Duration in Months (4)
Repayment ability Projected net cash flow plus projected inventory divided by a total line of credit (1), (2), (3)
Repayment history The average of loan principal repaid divided by principal due over the past three years (1), (3)
Tenure The ratio of acres owned and farmed to total acres farmed (3), (6)
Resident 1 if applicant is a resident; 0 otherwise (1),
Working experience Working experience in current profession in months (1),
Region and Country Region and Country of the Agricultural Business (1), (2), (4), (6)
Distance Distance to the microfinance institution in km (6)
Table 1: A List of Economic Metrics Applied in Literature
Developing a comprehensive agricultural credit scoring system requires careful consideration of the various metrics that influence creditworthiness. Below is a scheme for generating a composite score based on multiple metrics, including how to assign weights to each metric. We aim at eventually developing a comprehensive ESE scoring system by either obtaining real data or generating synthetic data according to this scheme in the future.
  1. 1.
    Identify and gather relevant metrics: Begin by identifying the metrics that are most critical to assessing creditworthiness in agricultural lending, such as past repayment history, crop yields, farm management practices, and market conditions.
  2. 2.
    Determine the relative importance of each metric: Assign weights to each metric based on its perceived importance in predicting creditworthiness. This can be done through a combination of expert judgment and statistical analysis of historical data. Weights may also be assigned based on the level of risk exposure associated with each ESE factor, with higher weights assigned to metrics that pose a greater risk to a farmer’s performance. Notable, the overall Thomson Reuters ESG Score is calculated by applying an automated, factual logic that determines the weight of each category. Each category consists of a different number of measures. The count of measures per category determines the weight of the respective category.
  3. 3.
    Normalize the metrics to ensure that each has the same scale, range, and units of measurement. This can be achieved through methods such as standardization or normalization.
  4. 4.
    Once the weights and normalized values of each metric have been determined, calculate the composite score by multiplying the weight of each metric by its normalized value and summing the products.
  5. 5.
    Validate the composite score by comparing it to actual default rates or repayment history. Refine the weights and metrics as needed based on the validation results.
  6. 6.
    Continuously monitor and update the credit scoring system by incorporating new data and adjusting weights and metrics as needed to improve the accuracy and predictive power of the model.
  7. 7.
    By following the above steps, a comprehensive agricultural credit scoring system can be developed that considers multiple metrics and assigns weights based on their relative importance in predicting creditworthiness.
Sector Metric How to Measure
E Soil health Measuring the quality and health of the soil, including factors such as nutrient levels, organic matter content, and erosion rates. Soil testing kits could be used to measure.
E Water Consumption Modelled evapotranspiration using climate data (adjusted for crop type and water availability). Water usage tracking software could be used as a tool to measure.
E Water Quality N & P concentration (measured in stream, or modeled using land cover for landscapes)
E Habitat Conversion Remotely sensed land use (of study area covered by natural habitat)
E Habitat Composition Remotely sensed land use (no. of land cover classes, indicating habitat diversity)
E Habitat Connectivity Remotely sensed land use (calculated connectivity score)
E Species Richness Field samples (richness relative to reference natural landscape)
E Species Abundance Field samples (abundance relative to reference natural landscape, i.e. the most undisturbed nearby similar habitat)
E Yield / Area Harvested mass of crop per unit area over a 5-year rolling average, perhaps using geometric mean to penalize variation in yield
E Use of pesticides and chemicals Liters of pesticide or chemicals used per hectare
E Carbon Footprint The carbon footprint for different crop types are shown in Sah and Devakumar (2018). Measuring the amount of greenhouse gas emissions generated from the farmer’s activities, including fertilizer use, transportation, and energy use.
S Health & safety issues Whether the applicant ever had safety issues while engaging in agricultural activities (0 if no, 1 if yes).
S Insuring health Whether the applicant has personal accident insurance or commercial insurance (0 if no, 1 if yes).
S Labor practices Measuring the fairness and safety of labor practices on the farm, including minimum wage compliance, safety regulations, and worker satisfaction.
S Community engagement Measuring the level of engagement and support the farmer provides to the local community, including economic benefits, charitable contributions, and community outreach.
S Ethical business practices Measuring the farmer’s adherence to ethical business practices, including transparency, honesty, and accountability.
S Compliance Measuring the farmer’s compliance with local and national regulations, including environmental, labor, and health and safety regulations.
Table 2: A List of ESG Metrics
3 A Framework for ESE-Joint Liability Model
The objective of the section is to examine a microcredit lending scenario in which a benevolent lender extends loans to a group of self-selected micro-entrepreneur borrowers who are jointly liable for repayment of the loan. Although each member of the group is given an individual loan, the group is collectively responsible for repaying the loans, and if the total loan is repaid, the entire group can qualify for a subsequent loan. The loans are then invested in a set of projects, which have an equal chance of success and can be either disjoint or correlated. It is assumed that the group members know the realized output of each project, while the lender does not. Both the lender and the borrowers have to make decisions in this scenario. The lender has to decide on the optimal contract structure, the degree of joint liability, and the level of punishment in the case of strategic default, while the borrowers have to decide whether to repay the loan or default.
One of the primary aims of this study is to determine the optimal group size that maximizes the benefit for the borrowers while ensuring that the lender breaks even. A larger group size can have a positive effect on the expected repayment amount, as more people are liable for repaying defaulted payments, thus ensuring a higher rate of repayment. However, a larger group size can also be a threat to repaying members, as they must repay all defaulting peers’ repayments when everyone else in the group defaults on repayment. Therefore, the section proposes a method to determine the optimal group size that enables joint liability to achieve its maximum loan ceiling. Another objective of this section is to derive the optimal Individual-ESE Score. The importance of a scoring system for both lenders and borrowers cannot be overstated. For lending companies, having access to a specific scoring mechanism for individual borrowers can provide a better understanding of their creditworthiness and repayment ability. This information is essential for insurance and lending companies in determining the risk associated with lending to a particular borrower. Additionally, the optimal group size of borrowers can be determined using the scoring system, enabling the company to mitigate risk by lending to a diversified portfolio. For borrowers, having a scoring system can give them a clearer picture of their ranking and eligibility for loans. This information is critical for farmers looking for group borrowing purposes as it can help them determine the number of peers needed to secure a loan and improve their chances of obtaining funding. Therefore, both lenders and borrowers can benefit from the implementation of a comprehensive scoring system.

### v2

@@ Abstract
This paper integrates a novel concept, coined as an individual–Environmental, Social, Economics (ESE) score with a joint liability model in order to gain insights into optimal group sizes and individual–ESE score relationships. We adopt a mean–variance utility function for farmers to effectively capture risk associated with anticipated profits. Through a set of simulation exercises, the paper studies a number of implications of incorporating $ESE$ scores into credit evaluation systems, offering, in the process, a nuanced comprehension of the repercussions under various climate conditions.
@@ 1 Introduction
In recent years the importance of incorporating sustainability factors into financial decision–making has gained a major traction. At the same time, while credit metrics for corporations have been extensively studied, there remains a noticeable gap in the assessment of individual borrowers, particularly farmers seeking credit to finance their agricultural operations. To address this gap, this paper introduces a novel concept of scoring system coined as an Environmental, Sustainability, and Economics (ESE), which considers not only traditional credit metrics, including economic factors, but importantly also environmental and social aspects which contribute to sustainable agricultural practices and which are so prominent in the standard Environmental, Sustainability, and Governance (ESG) score. See ztwp2024 for a further discussion and analyis of this individual–ESE concept.
@@
In particular, in this paper, we propose a novel link between an individual–ESE score and a joint liability model and provide lenders and borrowers (where lenders are insurance companies and borrowers are farmers) with optimal scores and corresponding optimal group sizes by maximizing total utility of a group of individual borrowers. We start our analysis with the simplest case of two borrowers in the group. Later we generalize it to a group of $n$ individual borrowers. It is important to incorporate the individual–ESE score mechanism into the joint liability model as this can benefit both lenders and borrowers. For lenders, this mechanism provides a better understanding of the borrower’s risk profile, enabling them to make informed lending decisions and, in the process, ensure more sustainable joint liability contracts. It also provides information about the optimal group size for lending purposes. For borrowers, the mechanism provides them with a clearer picture of their ranking relative to their peers, allowing them to identify strengths and weaknesses and work on improving their individual–ESE scores. It also provides information about the number of peers needed for a group borrowing purpose, helping borrowers form groups that are more likely to succeed and have improved access to credit. Therefore we argue that incorporating the individual–ESE score mechanism into credit evaluation processes can lead to better–informed lending decisions, improved access to credit, and sustainable development outcomes.
@@
The remaining parts of this paper are organized as follows. We introduce the set-up and assumptions of the model in Section 2. We derive a set of results from the models consisting of a group size of two and a group size of $n$ in Sections 3 and 4 respectively. In Section 5 we refine the utility functional form of the model to that of a mean variance, which incorporates both the expected value of return and risk simultaneously in the face of climate change. As the resulting model does not have clean and closed-form analytical expressions, a set of simulation exercises is provided based on a set of climate scenarios. Section 6 concludes the paper with several qualifications and avenues for future research.
2 The Framework for an Individual ESE-Joint Liability Model
We are interested in examining a microcredit lending scenario in which a benevolent lender extends loans to a group of self–selected micro–entrepreneur borrowers who are jointly liable for repayment of the loan. Although each member of the group is given an individual loan, the group is collectively responsible for repaying the loans, and if the total loan is repaid, the entire group can qualify for a subsequent loan. The loans are then invested in a set of projects, which have an equal chance of success and can be either disjoint or correlated. It is assumed that the group members know the realized output of each project, while the lender does not. Both the lender and the borrowers have to make decisions in this set up. The lender has to decide on the optimal contract structure, the degree of joint liability, and the level of punishment in the event of a strategic default, while the borrowers have to decide whether to repay or default the loan.
There are two prior studies which have inspired us in particular to write this paper. First rezaei2017optimal show that a group size can affect lending efficiency. Larger groups are more reliable in repaying loans for riskier projects, but the group size cannot become too large. Joint liability has a positive effect on borrowers’ repayment amount and welfare compared to individual liability, provided that group members impose strong punishment strategies against any strategically defaulting member. rezaei2017optimal also examine the possibility of relaxing some of the simplifying assumptions made in the basic model of joint liability. Second, in the context of climate change, castaing2021joint investigates the impact of risk–sharing networks on the adoption of weather shock management strategies in Burkina Faso; particularly, he studies how the principle of forced solidarity may reduce efforts to adopt techniques that mitigate exposure to climate change. The study finds that a system based on mutual assistance between farmers may reduce their incentives to adopt risk–mitigating strategies, both incrementally and transformationally. The study highlights the need for complementary mechanisms, such as insurance, to help farmers reduce their vulnerability to climate change.
This paper pursues two primary objectives. The first one is to determine an optimal group size that maximizes the benefit for the borrowers, while ensuring that the lender breaks even. A larger group size can have a positive effect on the expected repayment amount, as more people are liable for repaying defaulted payments, thus ensuring a higher rate of repayment. However, a larger group size can also pose a threat to repaying members, as they must repay all defaulting peers’ repayments, when everyone else in the group defaults on repayment. Therefore this paper proposes a method to determine the optimal group size that enables joint liability to achieve its maximum loan ceiling. A second objective of this paper is to derive an optimal individual–ESE score. The importance of such a scoring system for both lenders and borrowers cannot be overstated. For lending companies, having access to a specific scoring mechanism for individual borrowers can provide a better understanding of their creditworthiness and repayment prospect. This information is essential for lending companies in determining the risk associated with lending to a particular borrower. Additionally the optimal group size of borrowers can be determined using the scoring system, which enables the lending company to mitigate risk by lending to a diversified portfolio. For borrowers, having a scoring system can equip them with a clearer picture of their ranking and eligibility for loans. This information is critical for farmers as they search for a group for the borrowing purposes since this can help them determine the number of peers needed in a group to secure a loan and improve their chances of obtaining a funding. Therefore, both lenders and borrowers can benefit from the implementation of such a comprehensive scoring system.